A 10cm   inside diameter pipe carries a water flowing at  10 m/s.  Find its mass flow rate in kg/s  taking water density as   1 000  kg / cu m


1
Expert's answer
2021-11-17T10:37:49-0500

"A=\pi d^2/4=3.14\cdot0.1^2/4=0.00785\ (m^2)"


"Q=\rho Av=1000\cdot0.00785\cdot10=78.5\ (kg/s)" . Answer